OpenStudy (kyanda17):

Simplify (1/2)4.squared

OpenStudy (michele_laino):

is your expression like this: \[\Large {\left( {\frac{1}{2}} \right)^4}\]

OpenStudy (kyanda17):

yes

OpenStudy (michele_laino):

we can write this: \[\Large {\left( {\frac{1}{2}} \right)^4} = \frac{1}{{{2^4}}} = ...?\]

OpenStudy (mathstudent55):

Rule: \(\Large \left( \dfrac{a}{b} \right) ^n = \dfrac{a^n}{b^n} \) Apply the rule: \(\Large \left( \dfrac{1}{2} \right) ^4 = \dfrac{1^4}{2^4} = \dfrac{1 \times 1 \times 1 \times 1}{2 \times 2 \times 2 \times 2 } \) What is 1 * 1 * 1 * 1 = ? What is 2 * 2 * 2 * 2 = ?

OpenStudy (anonymous):

Would it not be (1/2) times (1/2) and 4 times 4?